Required information

[The following information applies to the questions displayed below.]

Delph Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year:

  

Molding Fabrication Total
Machine-hours 31,000 41,000 72,000
Fixed manufacturing overhead costs $ 710,000 $ 240,000 $ 950,000
Variable manufacturing overhead cost per machine-hour $ 5.80 $ 5.80

  

During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs:

  

Job D-70: Molding Fabrication Total
Direct materials cost $ 374,000 $ 321,000 $ 695,000
Direct labor cost $ 210,000 $ 130,000 $ 340,000
Machine-hours 21,000 10,000 31,000

  

Job C-200: Molding Fabrication Total
Direct materials cost $ 280,000 $ 260,000 $ 540,000
Direct labor cost $ 180,000 $ 280,000 $ 460,000
Machine-hours 10,000 31,000 41,000

Delph had no underapplied or overapplied manufacturing overhead during the year.

2. Assume Delph uses departmental predetermined overhead rates based on machine-hours.

a. Compute the departmental predetermined overhead rates.

b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200.

c. If Delph establishes bid prices that are 140% of total manufacturing costs, what bid prices would it have established for Job D-70 and Job C-200?

d. What is Delph’s cost of goods sold for the year?
Answer #1

Answers

  • All working forms part of the answer
  • Requirement 2 ‘a’

Working

Molding Department:

Fabrication Department:

A

Total Fixed Overheads

$              710,000

$             240,000

B

Total machine hours

                    31,000

                  41,000

C = A/B

Fixed overhead rate

$                   22.90

$                    5.85

D

Variable Overhead rate

$                     5.80

$                    5.80

E = C+D

Plantwide overhead rate

$                   28.70

$                 11.65

  • Requirement 2 ‘b’

Job D 70

Job C 200

Molding Department:

Direct material cost

$              374,000

$             280,000

Direct Labor cost

$              210,000

$             180,000

Manufacturing Overhead

$              602,700 [21000 x 28.70]

$             287,000 [10000 x 28.70]

Fabrication Department:

Direct material cost

$              321,000

$             260,000

Direct Labor cost

$              130,000

$             280,000

Manufacturing Overhead

$              116,500 [10000 x 11.65]

$             361,150 [31000 x 11.65]

Total manufacturing cost assigned

$          1,754,200

$         1,648,150

  • Requirement 2 ‘c’

Job D 70

Job C 200

A

Total manufacturing cost assigned

$        1,754,200

$               1,648,150

B = A x 140%

Bid Price

$        2,455,880

$               2,307,410

  • Requirement 2 ‘d’

--Since there were no beginning or ending work in process, cost of goods sold would be sum of cost of the two jobs completed and sold.

--Cost of Goods Sold = 1754200 + 1648150 = $ 3,402,350

Answer = $ 3,402,350